# MambaPEFT: Exploring Parameter-Efficient Fine-Tuning for Mamba

### v1

@@ 2.1 Mamba
$\overline{A}:=(I-d/2\cdot A)^{-1}(I+d/2\cdot A),\quad\overline{B}:=(I-d/2\cdot A)^{-1}dB,\quad\overline{C}:=C,\quad\overline{D}:=D$ (2)
@@
There have been numerous efforts to apply this mechanism to deep learning (e.g. convolutional representations (Gu et al., 2021b) and the initialization by Hippo matrices (Gu et al., 2020)). As a result, a time series model, Mamba (Gu & Dao, 2023), with a dynamic inference system called Selective State Space Model (S6), was developed as,
@@
Impressive results of Mamba in NLP have inspired researchers to adapt it to visual tasks. Vim (Zhu et al., 2024) is a pioneering work in this area. Similar to ViT (Dosovitskiy et al., 2020), it divides images into patches and then inputs the patch sequences into SSMs. There are other Mamba-based methods for visual tasks. For example, VMamba (Liu et al., 2024b) is a hybrid approach combining Mamba with 2D Convolutions. MambaVision (Hatamizadeh & Kautz, 2024) further integrates Attention. In this paper, we conduct experiments with Vim, which is a variant of the vanilla Mamba.
@@ 2.2 Parameter-Efficient Fine-Tuning for Transformer
Partial-tuning methods. While usual fine-tuning updates all parameters in the network, Partial-tuning updates a part of them. BitFit (Zaken et al., 2021) is a successful method in this category which tunes only the bias terms of the linear layers in Transformers. However, most of the linear layers in Mamba do not have the bias terms, which implies that it cannot be applied a priori.
@@
Reparameterization methods. Although reparameterization methods add low-rank adapters to Linear layers similar to the additive adapter methods, they adopt adapters that can be merged into the weights of the original Linear layer at the inference time. Due to this restriction, activation cannot be used in the adapter, while it eliminates the inference time overhead. The most notable one is LoRA (Hu et al., 2021), a method that reparameterizes the pre-trained Linear weights $W$ as $W^{\prime}=W+s\cdot W_{\mathrm{down}}\cdot W_{\mathrm{up}}$. While subsequent studies have improved the efficiency by finding ways to reparameterize the weights (Hayou et al., 2024; Liu et al., 2024a; Jiang et al., 2024), we first investigate how many ranks of the adapter should be attached to which linear weights in Mamba. In this regard, we will use a simple LoRA with few hyperparameters.
@@ 3.2 Re-designing and improving existing PEFT methods for Mamba
Prompt-tuning. Prompt-tuning can be applied directly to Mamba. As with Transformer, we simply add soft prompts to the input sequence. In normal Prompt-tuning, a prompt is inserted at the beginning of the input sequence. For Vim, however, we adjust the insertion position. This is because, unlike ViT, Mamba is a time-series model and works differently depending on where the prompt is inserted. In fact, Vim improved accuracy by inserting the class token in the middle of the sequence rather than the beginning. In addition, Mamba Register (Darcet et al., 2024) found it is preferable to insert the register tokens at equal intervals. Hence, in this paper, we construct three prompt types; prefix (resp. infix, suffix) type, inserting prompt tokens at the beginning (resp. with equal intervals, at the end).
@@ 3.4 Hybrid PEFT Search
In the first step, we restrict the search space to combinations of PEFT methods with minimal trainable parameters within each method, i.e., we explore whether to activate each PEFT method with fixed hyperparameters with minimal trainable parameters. This phase is based on our observation that performance degrades when too many PEFT methods or trainable parameters are added. In the second step, based on the PEFT method combination suggested by the first step, we greedily search the hyperparameters of both PEFT methods and their training. Since this phase inevitably increases the number of trainable parameters, we also include an option to remove a specific PEFT method. This approach helps to prevent excessive parameter growth and allows us to allocate more parameters to the more critical methods. We utilize the TPE algorithm (Bergstra et al., 2011) implemented in Optuna (Akiba et al., 2019) for the search in each step. Detailed algorithm is provided in Appendix B.
@@ 4 Experiments and discussion
Table 1: Test accuracy on the VTAB-1k benchmark pre-trained on ImageNet-1K using the training technique of DeiT (Touvron et al., 2021). The best hyperparameter settings maximizing the averaged accuracy of six datasets written in the ablation studies are used for each PEFT method. The number of trainable parameters are on the Cifar100 task. The best performance is highlighted in bold, and the second-best performance is underlined.
Model Method #Params (K) Natural Specialized Structured Avg.
@@
Vim-S Scratch 25,450 8.33 49.87 28.16 25.42
Full 25,450 59.35 68.74 34.39 47.08
Linear Probing 9 62.50 77.25 31.97 52.75
@@
Pos-embed-tuning 84 64.25 74.60 39.77 56.12
Bias-tuning 37 68.94 79.16 45.05 61.03
D-tuning 45 67.14 78.56 40.10 58.16
A-tuning 598 72.01 80.72 49.59 64.40
Conv1d-tuning 156 74.33 82.45 57.83 69.09
@@
Prompt-tuning 307 63.19 77.88 35.82 54.76
Affix-tuning (w/o proj) 230 63.90 77.66 34.21 54.30
Affix-tuning 117,000 75.84 83.29 58.94 70.29
Additional-scan 672 74.63 82.68 56.40 68.65
ParallelAdapter 663 76.10 83.97 59.97 70.96
@@
LoRA(x_proj) 2,540 74.41 81.92 54.88 67.77
LoRA(dt_proj) 2,442 75.35 83.05 57.12 69.30
LoRA(out_proj) 2,663 76.42 83.96 60.08 71.12
LoRA(in_proj) 1,483 76.58 84.08 60.16 71.25
LoRAp(d) 2,442 73.25 80.91 50.93 65.46
LoRAp(C) 2,417 72.78 81.57 51.35 65.61
LoRAp(B) 2,417 72.95 81.66 52.26 66.07
LoRAp(Z) 1,778 76.15 84.26 59.72 70.94
LoRAp(X) 1,778 76.64 83.89 60.84 71.52
@@
Hybrid (w/ proj) 117,236 77.00 84.41 61.55 72.05
Hybrid (w/o proj) 1,044 76.85 84.42 61.06 71.80
@@ 4.1 Experimental settings
Models. We use Vim-S (Zhu et al., 2024) as a base model in our experiments. We also experiment with ViT (Dosovitskiy et al., 2020) for comparison. We adopt pre-trained weights trained with ImageNet-1k (Deng et al., 2009) dataset using DeIT (Touvron et al., 2021) training framework in all models.
@@
Baselines. We adopt LoRA (Hu et al., 2021), Adaptformer (Chen et al., 2022), FacT-TK (Jie & Deng, 2023), and Adapter+ (Steitz & Roth, 2024) as existing PEFT methods for Transformers. For Mamba, we evaluate the 20 PEFT methods explained in Section 3. As a baseline for Hybrid PEFT, we use a combination of all PEFT methods in the experiments.
Implementation Details. We follow the setup of Jie & Deng (2023) in our experiments, using AdamW optimizer (Loshchilov & Hutter, 2017) and training the model for 100 epochs. The learning rate is set to 1e-3, with a cosine scheduler and a warmup period of 10 epochs. A weight decay with 1e-4 magnitude is applied. We do not perform data augmentation. For ablation studies, we use the average accuracy on six tasks: CIFAR-100, Sun397, Camelyon, Retinopathy, Clevr-Count, and sNORB-Elev.
@@ 4.2 Results for individual PEFT methods
Another notable finding regarding LoRA is that when it is attached to a linear weight with a small dimension, the accuracy continues to increase even if the rank of LoRA exceeds the full-rank, as shown in Figure 3. Full-rank refers to the smaller value of the input dimension or output dimension. This phenomenon is likely because the effective degrees of freedom are bounded by the full-rank. It follows that adding more parameters does not cause overfitting. Instead, it positively contributes to over-parameterization, similar to the findings in Ding et al. (2021). More details regarding the LoRA on large lineaer weights are described in Appendix Section A.1.
@@ 4.3 Results of Hybrid PEFT
Model Method #Params(%) BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c QBQA Avg.
Pythia 160M Full 100 58.7 62.2 39.0 29.1 52.1 42.3 23.9 28.4 42.0
@@ 5 Conclusion
The experiments demonstrate that Mamba benefits from PEFT more than Transformers. It improves accuracy without overfitting, even with a larger number of parameters, especially when fine-tuning data is limited. LoRA-based methods, particularly LoRAp(X), are effective with limited data. In contrast, the proposed Additional-scan method excels with larger datasets. We found that initialization significantly impacts performance when adding an additional dimension to A, despite Mamba’s robustness to initialization compared to other SSMs. Moreover, the proposed Affix-tuning was effective especially for large Mamba models. As to adapters on small linear weights, interestingly, over-parameterizing the original weight rather than decomposing it to a lower rank can be beneficial, which should holds for other architectures than Mamba. In hybrid PEFT, simply combining individually high-performance methods is inefficient. It is crucial to find the appropriate combination of PEFT methods. We believe that these results make a valuable contribution to the Mamba research community.
@@ A.1 Individual PEFT methods in detail
For Prompt-tuning, adding soft tokens to both the beginning and end is the most effective, as shown in Figure 4. Figure 4 also shows that the number of affixes improves performance up to 3 tokens. However, 4 or more affix tokens degrade performance, due to over-fitting caused by increasing the number of parameters.
A.2 Efficiency of PEFT methods for Mamba
@@ A.3 Per-task evaluation on VTAB-1k
We present the results for each task in VTAB-1k, including different base model sizes and training data. Specifically, we also report the accuracy on other base model architectures such as Vim-tiny (Zhu et al., 2024), ViT-tiny (Touvron et al., 2021), and ViT-B (Dosovitskiy et al., 2020), as well as different pre-training datasets such as ImageNet-21k (Deng et al., 2009) or using the training framework of (Dosovitskiy et al., 2020). The compared PEFT methods for Transformers are LoRA (Hu et al., 2021), Adaptformer (Chen et al., 2022), Adapter+ (Steitz & Roth, 2024), FacT-TK (Jie & Deng, 2023), Head2Toe (Evci et al., 2022), VQT (Tu et al., 2023), BitFit (Zaken et al., 2021), and LoSA (Mercea et al., 2024). The results are shown in Section A.3.
@@
Full 68.9 87.7 64.3 97.2 86.9 87.4 38.8 75.9 79.7 95.7 84.2 73.9 83.4 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 47.6 68.9
Linear 64.4 85.0 63.2 97.0 86.3 36.6 51.0 69.1 78.5 87.5 68.5 74.0 77.1 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 26.9 57.6
BitFit 72.8 87.0 59.2 97.5 85.3 59.9 51.4 73.3 78.7 91.6 72.9 69.8 78.3 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 44.1 65.2
LoRA 67.1 91.4 69.4 98.8 90.4 85.3 54.0 79.5 84.9 95.3 84.4 73.6 84.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0 59.8 74.5
FacT-TK 70.6 90.6 70.8 99.1 90.7 88.6 54.1 80.6 84.8 96.2 84.5 75.7 85.3 82.6 68.2 49.8 80.7 80.8 47.4 33.2 43.0 60.7 75.6
Adapter+ 83.7 94.2 71.5 99.3 90.6 88.2 55.8 83.3 87.5 97.0 87.4 72.9 86.2 82.9 60.9 53.7 80.8 88.4 55.2 37.3 46.9 63.3 77.6
LoSA 82.5 92.8 76.1 99.7 90.5 82.0 55.8 82.8 86.6 97.1 87.0 76.7 86.9 81.5 62.3 48.6 82.1 94.2 61.7 47.9 45.6 65.5 78.4

### v2

@@ 2.1 Mamba
$\overline{A}:=\exp(dA),\quad\overline{B}:=(dA)^{-1}(\exp(dA)-I)dB,\quad\overline{C}:=C,\quad\overline{D}:=D$ (2)
@@
While many SSM-based DNN have tuned the A, B, C, and D (Gu et al., 2020; 2021b), Selective State Space Model (S6) was introduced in Mamba (Gu & Dao, 2023), in which the A, B, C, and D are dynamic and data-dependent as follows:
@@
Impressive results of Mamba in NLP have inspired researchers to adapt it to visual tasks. Vim (Zhu et al., 2024) is a pioneering work in this area. Similar to ViT (Dosovitskiy et al., 2020), it divides images into patches and then inputs the patch sequences into SSMs. Vim improved performance in visual tasks by processing patches bidirectionally with S6 and placing a class token in the middle of the patch tokens instead of at the beginning. There are other Mamba-based methods for visual tasks. For example, VMamba (Liu et al., 2024b) is a hybrid approach combining Mamba with 2D Convolutions. MambaVision (Hatamizadeh & Kautz, 2024) further integrates Attention. In this paper, we conduct experiments with Vim, which is a variant of the vanilla Mamba.
@@ 2.2 Parameter-Efficient Fine-Tuning for Transformer
Partial-tuning methods. While usual fine-tuning updates all parameters in the network, Partial-tuning updates a part of them. BitFit (Zaken et al., 2021) is a successful method in this category which tunes only the bias terms of the linear layers in Transformers. However, most of the linear layers in Mamba do not have the bias terms, which implies that it cannot be applied a priori. Other methods use pruning to decide where to tune Zhang et al. (2024c); He et al. (2023).
@@
Reparameterization methods. Although reparameterization methods add low-rank adapters to Linear layers similar to the additive adapter methods, they adopt adapters that can be merged into the weights of the original Linear layer at the inference time. Due to this restriction, activation cannot be used in the adapter, while it eliminates the inference time overhead. The most notable one is LoRA (Hu et al., 2021), a method that reparameterizes the pre-trained Linear weights $W$ as $W^{\prime}=W+s\cdot W_{\mathrm{down}}\cdot W_{\mathrm{up}}$. While subsequent studies have improved the efficiency by finding ways to reparameterize the weights (Lian et al., 2022; Hayou et al., 2024; Liu et al., 2024a; Jiang et al., 2024), we first investigate how many ranks of the adapter should be attached to which linear weights in Mamba. In this regard, we will use a simple LoRA with few hyperparameters.
@@ 3.2 Re-designing and improving existing PEFT methods for Mamba
Prompt-tuning. Prompt-tuning can be applied directly to Mamba. As with Transformer, we simply add soft prompts to the input sequence. In normal Prompt-tuning, a prompt is inserted at the beginning of the input sequence. For Vim, however, we adjust the insertion position. This is because, unlike ViT, Mamba is a time-series model and works differently depending on where the prompt is inserted. In fact, Vim improved accuracy by inserting the class token in the middle of the sequence. In addition, Mamba Register (Darcet et al., 2024) found it is preferable to insert the register tokens at equal intervals. Hence, in this paper, we construct three prompt types; prefix (resp. infix, suffix) type, inserting prompt tokens at the beginning (resp. with equal intervals, at the end).
@@ 3.4 Hybrid PEFT Search
In the first step, we restrict the search space to combinations of PEFT methods with minimal trainable parameters within each method, i.e., we explore whether to activate each PEFT method with fixed hyperparameters with minimal trainable parameters. This phase is based on our observation that performance degrades when too many PEFT methods or trainable parameters are added. In the second step, based on the PEFT method combination suggested by the first step, we greedily search the hyperparameters of both PEFT methods and their training. Since this phase inevitably increases the number of trainable parameters, we also include an option to remove a specific PEFT method. This approach helps to prevent excessive parameter growth and allows us to allocate more parameters to the more critical methods. We utilize the TPE algorithm (Bergstra et al., 2011) implemented in Optuna (Akiba et al., 2019) for the search in each step. The detailed algorithm is provided in Appendix B.
@@ 4 Experiments and discussion
Table 1: Test accuracy on the VTAB-1k benchmark pre-trained on ImageNet-1K using the training technique of DeiT (Touvron et al., 2021). The best hyperparameter settings maximizing the averaged accuracy of six datasets written in the ablation studies are used for each PEFT method. The Time Ratio represents the ratio of the training time to the time taken for full fine-tuning. The number of trainable parameters is the average of all tasks.
Model Method #Params (K) Time Ratio Natural Specialized Structured Avg.
@@
Vim-S Scratch 25,450 1.00 8.33 49.87 28.16 25.42
Full 25,450 1.00 59.35 68.74 34.39 47.08
Linear Probing 9 0.19 62.50 77.25 31.97 52.75
@@
Pos-embed-tuning 84 0.52 64.25 74.60 39.77 56.12
Bias-tuning 37 0.54 68.94 79.16 45.05 61.03
D-tuning 45 0.56 67.14 78.56 40.10 58.16
A-tuning 598 0.54 72.01 80.72 49.59 64.40
Conv1d-tuning 156 0.54 74.33 82.45 57.83 69.09
@@
Prompt-tuning 307 0.54 69.92 79.20 47.75 62.54
Affix-tuning (w/o proj) 230 0.59 72.26 81.03 50.73 65.04
Affix-tuning 117,000 0.66 75.84 83.29 58.94 70.29
Additional-scan 672 0.65 74.63 82.68 56.40 68.65
ParallelAdapter 663 0.55 76.10 83.97 59.97 70.96
@@
LoRA(x_proj) 2,540 0.57 74.41 81.92 54.88 67.77
LoRA(dt_proj) 2,442 0.57 75.35 83.05 57.12 69.30
LoRA(out_proj) 2,663 0.57 76.42 83.96 60.08 71.12
LoRA(in_proj) 1,483 0.67 76.58 84.08 60.16 71.25
LoRAp(d) 2,442 0.63 73.25 80.91 50.93 65.46
LoRAp(C) 2,417 0.57 72.78 81.57 51.35 65.61
LoRAp(B) 2,417 0.65 72.95 81.66 52.26 66.07
LoRAp(Z) 1,778 0.67 76.15 84.26 59.72 70.94
LoRAp(X) 1,778 0.66 76.64 83.89 60.84 71.52
@@
All LoRA (r=8) 1,228 0.80 76.11 84.32 59.92 71.02
LoRA(in_proj+out_proj) 709 0.72 75.69 84.42 59.43 70.68
Hybrid (w/ proj) 117,236 0.86 77.00 84.41 61.55 72.05
Hybrid (w/o proj) 1,044 0.83 76.85 84.42 61.06 71.80
@@ 4.1 Experimental settings
Models. We use Vim-S (Zhu et al., 2024) as a base model in our experiments. We also experiment with ViT (Dosovitskiy et al., 2020) for comparison. We adopt pre-trained weights trained with ImageNet-1k (Deng et al., 2009) using DeIT (Touvron et al., 2021) training framework in all models.
@@
Baselines. We adopt LoRA (Hu et al., 2021), Adaptformer (Chen et al., 2022), FacT-TK (Jie & Deng, 2023), SPT-LoRA (He et al., 2023), and Adapter+ (Steitz & Roth, 2024) as existing PEFT methods for Transformers. For Vim, we evaluate the 20 PEFT methods explained in Section 3. As a baseline for Hybrid PEFT, we use a combination of all PEFT methods in the experiments.
Implementation Details. We follow the setup of Jie & Deng (2023) in our experiments, using AdamW optimizer (Loshchilov & Hutter, 2017) and training the model for 100 epochs. The learning rate is set to 1e-3, with a cosine scheduler and a warmup period of 10 epochs. A weight decay with 1e-4 magnitude is applied. We do not perform data augmentation. For ablation studies, we use the average accuracy on six tasks: CIFAR-100, Sun397, Camelyon, Retinopathy, Clevr-Count, and sNORB-Elev. In several previous studies, hyperparameters were tuned for each task, and the maximum validation accuracy during training was reported. In contrast, we tuned the hyperparameters to maximize the average of six tasks used in ablation, ensuring that the results obtained are universal and beneficial for many future studies. Additionally, the evaluation is conducted using the weights from the final epoch.
@@ 4.2 Results for individual PEFT methods
Another notable finding regarding LoRA is that when it is attached to a linear weight with a small dimension, the accuracy continues to increase even if the rank of LoRA exceeds the full-rank, as shown in Figure 3. Full-rank refers to the smaller value of the input dimension or output dimension. This phenomenon is likely because the effective degrees of freedom are bounded by the full-rank. It follows that adding more parameters does not cause overfitting. Instead, it positively contributes to over-parameterization, similar to the findings in Ding et al. (2021). More details regarding the LoRA on large linear weights are described in Appendix Section A.1.
@@ 4.3 Results of Hybrid PEFT
Model Method #Params(%) BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
Pythia 160M Full 100 61.3 62.9 37.1 30.7 50.6 41.5 24.3 27.8 42.0
@@ 5 Conclusion
Our findings are as follows. (1) Mamba benefits from PEFT more than Transformers. It improves accuracy without overfitting, even with a larger number of parameters in PEFT (Compare Figure 4 and Figure 4). (2) LoRA-based methods, particularly LoRAp(X), are effective with limited data (See Table 1). (3) In contrast, the proposed Additional-scan excels with larger datasets (See Table 3 and Figure 7). (4) We found that initialization significantly impacts performance when adding an additional dimension to A, despite Mamba’s robustness to initialization compared to other SSMs (See Figure 3). (4) The proposed Affix-tuning was effective, especially for large Mamba models (See Table 3). (5) When adding LoRA on small linear weights, interestingly, over-parameterizing the original weight rather than decomposing it to a lower rank can be beneficial, which should hold for other architectures than Mamba (Figure 3). (6) In hybrid PEFT, simply combining individually high-performance methods, as many previous works did, is inefficient. It is crucial to find the appropriate combination of PEFT methods. (See Table 1 and Table 8) We believe that these results make a valuable contribution to the Mamba research community.
@@ A.1 Individual PEFT methods in detail
For Prompt-tuning, adding soft tokens to both the beginning and end is the most effective, as shown in Figure 4. Figure 4 also shows that the number of affixes improves performance up to 3 tokens. However, 4 or more affix tokens degrade performance due to over-fitting caused by increasing the number of parameters.
A.2 Efficiency of PEFT methods for Vision Mamba
@@ A.3 Per-task evaluation on VTAB-1k
We present the results for each task in VTAB-1k, including different base model sizes and training data. Specifically, we also report the accuracy on other base model architectures such as Vim-B, Vim-tiny (Zhu et al., 2024), ViT-tiny (Touvron et al., 2021), and ViT-B (Dosovitskiy et al., 2020), as well as different pre-training datasets such as ImageNet-21k (Deng et al., 2009) or using the training framework of (Dosovitskiy et al., 2020). The information of Vim-B is recently released, so we additionally experimented the performance here in the Appendix. The compared PEFT methods for Transformers are LoRA (Hu et al., 2021), Adaptformer (Chen et al., 2022), Adapter+ (Steitz & Roth, 2024), FacT-TK (Jie & Deng, 2023), Head2Toe (Evci et al., 2022), VQT (Tu et al., 2023), BitFit (Zaken et al., 2021), SPT-LoRA (He et al., 2023), and LoSA (Mercea et al., 2024). The results are shown in Section A.3.
@@
Full 68.9 87.7 64.3 97.2 86.9 87.4 38.8 75.9 79.7 95.7 84.2 73.9 83.4 56.3 58.6 41.7 65.5 57.5 46.7 25.7 29.1 47.6 65.6
Linear 64.4 85.0 63.2 97.0 86.3 36.6 51.0 69.1 78.5 87.5 68.5 74.0 77.1 34.3 30.6 33.2 55.4 12.5 20.0 9.6 19.2 26.9 52.3
BitFit 72.8 87.0 59.2 97.5 85.3 59.9 51.4 73.3 78.7 91.6 72.9 69.8 78.3 61.5 55.6 32.4 55.9 66.6 40.0 15.7 25.1 44.1 62.1
LoRA 67.1 91.4 69.4 98.8 90.4 85.3 54.0 79.5 84.9 95.3 84.4 73.6 84.6 82.9 69.2 49.8 78.5 75.7 47.1 31.0 44.0 59.8 72.3
FacT-TK 70.6 90.6 70.8 99.1 90.7 88.6 54.1 80.6 84.8 96.2 84.5 75.7 85.3 82.6 68.2 49.8 80.7 80.8 47.4 33.2 43.0 60.7 73.2
Adapter+ 83.7 94.2 71.5 99.3 90.6 88.2 55.8 83.3 87.5 97.0 87.4 72.9 86.2 82.9 60.9 53.7 80.8 88.4 55.2 37.3 46.9 63.3 75.5
LoSA 82.5 92.8 76.1 99.7 90.5 82.0 55.8 82.8 86.6 97.1 87.0 76.7 86.9 81.5 62.3 48.6 82.1 94.2 61.7 47.9 45.6 65.5 76.4
@@
SPT-LoRA 62.1 89.8 69.7 92.6 89.0 91.6 43.4 76.9 85.6 95.9 85.6 73.8 85.2 76.3 55.8 49.2 82.1 78.3 55.4 31.8 37.1 58.3 70.8
Adapter+ 61.6 91.0 68.6 92.5 91.8 89.8 43.2 76.9 85.5 95.1 85.3 73.7 84.9 78.4 58.8 50.7 81.2 83.6 55.1 32.7 35.6 59.5 71.3
[…]
Full 28.5 65.0 5.5 63.2 42.4 86.4 16.5 43.9 65.1 83.5 55.8 73.6 69.5 22.0 44.1 20.7 33.3 36.1 44.2 7.1 14.9 27.8 42.5
Linear 46.4 86.8 63.9 80.0 90.2 43.1 41.3 64.5 82.2 89.5 71.4 73.8 79.2 37.7 38.0 36.3 63.6 16.8 26.7 16.2 22.6 32.2 54.0
Affix-tuning (w/o proj) 59.0 88.1 66.8 84.5 90.8 52.5 43.0 69.2 79.3 93.5 82.2 73.8 82.2 62.0 54.3 43.5 76.8 59.9 50.7 25.2 33.8 50.8 64.2
Affix-tuning 62.2 88.8 68.3 91.2 91.3 88.9 43.7 76.4 83.0 94.1 83.5 73.5 83.5 76.4 59.1 33.7 79.2 79.6 50.8 31.1 35.3 55.6 69.1
Additional-scan 60.3 88.3 68.9 90.2 91.7 85.7 43.2 75.5 83.0 93.7 82.1 70.7 82.4 74.1 61.7 44.3 80.2 73.5 46.8 29.4 36.8 55.8 68.7
ParallelAdapter 62.8 89.8 68.4 91.5 91.2 88.2 44.0 76.6 83.5 95.0 85.0 72.2 83.9 83.3 64.2 50.2 79.3 78.9 53.0 33.8 40.6 60.4 71.3
LoRA(out_proj) 63.8 89.5 68.0 91.7 92.1 87.7 44.2 76.7 83.9 94.8 85.0 72.4 84.0 83.6 63.4 50.3 81.3 76.3 52.5 33.0 40.6 60.1 71.3
LoRA(in_proj) 63.7 90.2 67.7 91.0 90.8 89.7 43.9 76.7 85.0 95.1 84.6 72.5 84.3 81.3 60.8 50.8 79.9 82.2 50.9 31.4 37.6 59.4 71.0
LoRAp(X) 63.9 90.1 68.4 91.2 91.0 90.4 43.9 77.0 84.1 95.2 84.0 72.5 83.9 83.0 62.7 49.6 79.9 81.7 51.4 31.2 39.9 59.9 71.3